OpenStudy (anonymous):

First, solve the equation for y. Second, find the value of y for the given x-value. 3x - 2y = -7, when x = 3 Please help! And fast!

OpenStudy (opcode):

\[3x - 2y = -7\] \[3 x+7 = 2 y\] \[y = \dfrac{(3 x)+7}{2}\] \[x = 3\] \[y = \dfrac{3 (3)+7}{2}\] \[y = \dfrac{9+7}{2}\] \[y = \dfrac{16}{2}\] \[y = 8\]

OpenStudy (anonymous):

Thank you all! :)